# FLAME: Learning to Navigate with Multimodal LLM in Urban Environments

### v1

@@ 4.1 Experiment Setup
For the original VLN datasets, we employ three metrics for performance evaluation: Task Completion (TC), Shortest-Path Distance (SPD) and Normalized Dynamic Time Warping (nDTW). Specifically, TC represents the percentage of successful navigation. SPD calculates the minimum distance from the stop location to the goal. nDTW assesses the overlap between the agent’s and the ground truth trajectories.
@@
Our agent is built upon Otter and OpenFlamingo (Li et al. 2023a; Awadalla et al. 2023), integrating CLIP (Radford et al. 2021) and LLaMA (Touvron et al. 2023), trained on a single A100 GPU. To accommodate CLIP’s input size, we center-crop and resize panoramas, which differs from the broader visual context used in existing models but aligns better with MLLM’s nature.
@@
4.2 Comparison with SOTAs
In the section, we compare FLAME with previous state-of-the-art (SOTA) approaches. As shown in Table 1, FLAME establishes new SOTA performance on both Touchdown and Map2seq datasets. On Touchdown test split, FLAME surpasses the previous SOTA, Loc4Plan (Tian et al. 2024), by 7.3% in TC and %1.97 in SPD. This demonstrates the superiority of our MLLM-based approach in comprehending navigational instructions and environmental cues, resulting in higher success rates and better path adherence. For Map2seq, FLAME outperforms VELMA (Schumann et al. 2023), an LLM-based method, with a 3.74% increase in TC and a 5.35% improvement in nDTW. This highlights the advantage of Multimodal LLMs in capturing comprehensive information compared to text-only LLM, which may introduce information loss when translating visual data.
Furthermore, we evaluated open-sourced methods in our visual setting, which features a restricted field of view similar to human vision. The results, marked with an asterisk (*) in Table 1, show a significant performance drop for baseline models, especially on the Touchdown dataset. This underscores the dependence of baseline methods on panoramic environmental perception. In contrast, FLAME excels at navigating with limited visual input, more closely approximating human navigation capabilities.
4.3 Reasoning Performance
We evaluated FLAME’s reasoning capabilities during navigation using the self-consistency approach (Wang et al. 2023b), exploring various decoding paths and temperatures. The results are presented in Table 2. Rationale Coherence (RC) and Rationale-Action Alignment (RA) consistently remained above 80% and 95% respectively, indicating robust rationale generation and strong consistency. Higher temperatures led to performance fluctuations, especially with fewer decoding paths. However, when we increase the number of decoding paths to 8, we see more pronounced improvements in both TC and RC. On Touchdown, FLAME achieved optimal TC performance with a temperature of 1.0 and 8 decoding paths, outperforming greedy decoding (T=0.0, P=1) by 1.72% in TC and 1.75% in RC. For Map2seq, with 8 decoding paths, FLAME surpassed greedy decoding by 3.91% in TC and 3.62% in RC on the test set. These results demonstrate that increased sampling diversity and a larger decoding budget enable FLAME to generate diverse rationales and effectively ensemble reasoning results, leading to improved decision-making. This provides strong evidence for the reasoning capabilities developed through synthetic rationale tuning and the advanced decision-making capacity of FLAME’s architecture.
@@ 4.3 Reasoning Performance
To validate the reliability of our automatic metric calculations (Eqs. 7 and 8), we conducted human evaluations on 50 instances each from the Touchdown and Map2seq datasets. The results, presented in Table 3, revealed discrepancies between the vanilla method and human assessments. However, our calibrated approach significantly reduced these disparities, demonstrating the human-comparable reliability of the automatic evaluation method. Consequently, we employ this calibrated evaluation technique as default. Specifically, We calibrate the outputs of $\text{CFR}_{\text{rc}}$ and $\text{CFR}_{\text{ra}}$ as follows:
@@ 4.4 Analyses
Table 4 presents our investigation into the impact of three-phase tuning on navigation performance. The first row, where both the first phase (P1) and second phase (P2) are omitted, demonstrates sub-optimal navigation performance. Implementing the first phase tuning yields increases of 0.99% and 1.12% in TC for dev splits of Touchdown and Map2seq, respectively, suggesting improved environmental familiarity. The second phase of tuning leads to optimal performance across both datasets, indicating the importance of tuning with sequential observations. These results underscore the crucial role of phased learning in equipping the MLLM with advanced navigational skills, from single perception to sequential trajectory understanding, facilitating adaptation from general scenarios. Notably, even the non-phased training surpasses current state-of-the-art results, further validating the efficacy of FLAME’s architecture.

### v2

@@ 4 Experiments
For the original VLN datasets, we employ three metrics for performance evaluation: Task Completion (TC), Shortest-Path Distance (SPD) and Normalized Dynamic Time Warping (nDTW). Specifically, TC represents the percentage of success. SPD calculates the minimum distance from the stop location to the goal. nDTW assesses the overlap between the agent’s and the ground truth trajectories.
@@
Our agent is built upon Otter and OpenFlamingo (Li et al. 2023a; Awadalla et al. 2023), integrating CLIP (Radford et al. 2021) and LLaMA (Touvron et al. 2023). To accommodate CLIP’s input size, we crop and resize panoramas, which differs from the broader visual context used in existing models but aligns better with MLLM’s nature. In the Touchdown task, we randomize the agent’s heading at the start of each trajectory by selecting one of the possible directions based on its neighboring nodes. Training for the first two phases takes 1 hour each, while the navigation fine-tuning requires 12 hours on a single A100 GPU.
@@
Comparison with SOTAs
In the section, we compare FLAME with previous state-of-the-art (SOTA) approaches. As shown in Table 1, FLAME establishes new SOTA performance on both Touchdown and Map2seq datasets. On Touchdown test split, FLAME surpasses the previous SOTA, Loc4Plan (Tian et al. 2024), by 7.3% in TC and 1.97% in SPD. This demonstrates the superiority of our MLLM-based approach in comprehending navigational instructions and environmental cues, resulting in higher success rates and better path adherence. For Map2seq, FLAME outperforms VELMA (Schumann et al. 2024), an LLM-based method, with a 3.74% increase in TC and a 5.35% improvement in nDTW. This highlights the advantage of Multimodal LLMs in capturing comprehensive information compared to text-only LLM.
Furthermore, we evaluated open-sourced methods in our visual setting, which features a restricted field of view similar to human vision. The results, marked with an asterisk (*) in Table 1, show a significant performance drop for baseline models, especially on the Touchdown dataset. This underscores the dependence of baseline methods on panoramic environmental perception. In contrast, FLAME excels at navigating with non-panoramic visual input.
Reasoning Performance
We evaluated FLAME’s reasoning capabilities during navigation using the self-consistency approach (Wang et al. 2023), exploring various decoding paths and temperatures. The results are presented in Table 2. Rationale Coherence (RC) and Rationale-Action Alignment (RA) consistently remained above 80% and 95% respectively, indicating robust rationale generation and strong consistency. Higher temperatures led to performance fluctuations, especially with fewer decoding paths. However, when we increase the number of decoding paths to 8, we see more pronounced improvements in both TC and RC. On Touchdown, FLAME achieved optimal TC performance with a temperature of 1.0 and 8 decoding paths, outperforming greedy decoding (T=0.0, P=1) by 1.72% in TC and 1.75% in RC. For Map2seq, with 8 decoding paths, FLAME surpassed greedy decoding by 3.91% in TC and 3.62% in RC on the test set. These results demonstrate that increased sampling diversity and a larger decoding budget enable FLAME to generate diverse rationales and effectively ensemble reasoning results, leading to improved decision-making. This provides strong evidence for the reasoning capabilities developed through synthetic rationale tuning and the advanced decision-making capacity of FLAME’s architecture.
@@
To validate the reliability of our automatic metric calculations (Eqs. 9 and 10), we conducted human evaluations on 50 instances each from the Touchdown and Map2seq datasets. The results, presented in Table 3, revealed discrepancies between the vanilla method and human assessments. However, our calibrated approach significantly reduced these disparities, demonstrating the human-comparable reliability of the automatic evaluation method. Consequently, we employ this calibrated evaluation technique as default. Specifically, We calibrate the outputs of $\text{CFR}_{\text{rc}}$ and $\text{CFR}_{\text{ra}}$ as follows:
@@
Table 4 presents our investigation into the impact of three-phase tuning on navigation performance. The first row, which corresponds to the vanilla MLLM checkpoint that bypassed the first (P1) and second (P2) phases, demonstrates sub-optimal navigation performance. Implementing the first phase tuning yields increases of 0.99% and 1.12% in TC for dev splits of Touchdown and Map2seq, respectively, suggesting improved environmental familiarity. The second phase of tuning leads to optimal performance across both datasets, indicating the importance of tuning with sequential observations. These results underscore the crucial role of phased learning in equipping the MLLM with advanced navigational skills, from single perception to sequential trajectory understanding, facilitating adaptation from general scenarios. Notably, even the non-phased training surpasses current state-of-the-art results, further validating the efficacy of FLAME’s architecture.
@@ 5 Conclusion
Acknowledgments
This work was supported in part by the Natural Science Foundation of China under Grant 62303307, in part by National Key R&D Program of China under Grant No.2023YFB4705700, in part by Shanghai Municipal Science and Technology Major Project under Grant 2021SHZDZX0102 and in part by the Fundamental Research Funds for the Central Universities.